# Towards Measuring and Modeling “Culture” in LLMs: A Survey

### v3

@@ Abstract
We present a survey of 55 recent papers that aim to study cultural representation and inclusion in large language models. We observe that none of the studies define “culture,” which is a complex, multifaceted concept; instead, they probe the models on some specially designed datasets which represent certain aspects of “culture.” We call these aspects the proxies of cultures, and organize them across three dimensions of demographic, semantic and linguistic-cultural interaction proxies. We also categorize the probing methods employed. Our analysis indicates that only certain aspects of “culture,” such as values and objectives, have been studied, leaving several other interesting and important facets, especially the multitude of semantic domains Thompson et al. (2020) and aboutness Hershcovich et al. (2022), unexplored. Two other crucial gaps are the lack of robustness and situatedness of the current methods. Based on these observations, we provide several recommendations for a holistic and practically useful research agenda for furthering cultural inclusion in LLMs and LLM-based applications. Compilation and details of papers used for the survey can be found via our GitHub repository [^1] https://github.com/faridlazuarda/cultural-llm-papers
@@
Muhammad Farid Adilazuarda1∗, Sagnik Mukherjee1thanks: Equal contribution, Pradhyumna Lavania2, Siddhant Singh2, Ashutosh Dwivedi2, Alham Fikri Aji1, Jacki O’Neill3, Ashutosh Modi2, Monojit Choudhury1 1MBZUAI  2Indian Institute of Technology Kanpur, India 3Microsoft Africa Research Institute, Kenya farid.adilazuarda@mbzuai.ac.ae, mukherjeesagnik2314@gmail.com
@@ 1 Introduction
Recently, there have been several studies on socio-cultural aspects of Large Language Models (LLMs) spanning from aligning them to human values and safety measures Glaese et al. (2022); Bai et al. (2022b, a) to studying LLMs as personas belonging to certain cultures Gupta et al. (2024); Kovač et al. (2023) and even their skills for resolving moral dilemmas in the context of value pluralistic considerations Sorensen et al. (2023); Tanmay et al. (2023). In order to make LLMs inclusive and deployable across regions and applications, it is indeed necessary for these models to be able to function adequately under different “cultural” contexts. The growing body of work that broadly aims at evaluating LLMs for their multi-cultural awareness and biases underscore an important problem - that the existing models are strongly biased towards Western, Anglo-centric or American “cultures” Johnson et al. (2022); EVS/WVS (2022); Hofstede (1984); Cieciuch and Schwartz (2012); Dwivedi et al. (2023). Such biases are arguably detrimental to the performance of the models in non-Western contexts leading to disparate utility, potential for unfairness across regions, and concerns about the impact on global cultural diversity, since if biased models reinforce dominant cultures, whether implicitly or explicitly, they might lead to a cycle of cultural homogeneity Vaccino-Salvadore (2023); Schramowski et al. (2021). The most recent generation of LLMs, with their impressive ability and widespread availability, only make this issue more pressing. It is therefore a timely moment to review the literature on LLMs and culture.
@@
In this work, we survey 39 NLP papers that study cultural representation, awareness or bias in LLMs either explicitly (e.g., Huang and Yang (2023a), Zhou et al. (2023), Cao et al. (2024b)) or implicitly (e.g., Wan et al. (2023)). It is quickly apparent that these papers either do not attempt to define culture or where they do, use very high-level definitions. For example, a common definition is “the way of life of a collective group of people, [that] distinguishes them from other groups with other cultures” (Mora, 2013; Shweder et al., 2007; Hershcovich et al., 2022). Not only do the papers typically use broad-brush definitions, most do not engage in a critical discussion on the topic. [^2] The situation is similar to that described in  Blodgett et al. (2020) in the context of research on “bias”. This is perhaps unsurprising as “culture” is a concept which evades simple definition. It is multifaceted, meaning different things to different people at different times. For example, some of the many and often implicitly applied meanings of culture include a.) “Cultural Heritage” such as art, music, and food habits [^3] https://uis.unesco.org/sites/default/files/documents/analysis_sdg_11.4.1_2022_final_alt_cover_0.pdf Blake (2000), b.) “Interpersonal Interactions” between people from different backgrounds (e.g., ways of speaking in a meeting, politeness norms) Monaghan et al. (2012), or c.) The “Ways of Life” of a collective group of people distinguishing them from other groups. There are a variety of sociological descriptions of culture, for example, Parsons (1972) describes it as the the pattern of ideas and principles which abstractly specify how people should behave, but which do so in ways which prove practically effective relative to what people want to do Münch et al. (1992); Parsons (1972). However, these too are high-level and hard to concretise. Further complications arise because the instantiation of culture is necessarily situated. Every individual and group lies at the intersection of multiple cultures (defined by their political, professional, religious, regional, class-based and other affiliations) and these are invoked according to the situation, typically in contrast to another group(s).
@@
How then is culture handled in NLP research? As we shall demonstrate, the datasets and studies are typically designed to tease out the differential performance of the models across some set of variables. Whilst often not addressed explicitly, it is in the very choice of datasets that authors specify the features of culture they will examine. That is, the datasets themselves can be considered to be proxies for culture. As such it is useful to examine the different instantiations of culture found in this research. From food and drink, to norms and values, how have researchers represented culture in and through their datasets? In doing so we attempt to make explicit the various facets of culture which have been studied, and to highlight gaps in the research. We call for a more explicit acknowledgment of the link between the datasets employed and the facets of culture studied, and hope that the schema described in this paper provides a useful mechanism for this. In addition, we highlight limitations in the robustness of the probing methods used in the studies, which raises doubts about the reliability and generalizability of the findings.
@@
Finally, we believe this literature survey highlights the limitations of NLP benchmarking for understanding how a concept as broad and complex as culture plays out in LLMs. Whilst benchmarking is important and necessary, it is not sufficient, as the choices made in creating rigorous benchmarking datasets are unlikely to reveal the full extent of either LLMs cultural limitations or their full cultural representation. Not only is culture multi-faceted, but cultural representation is tied in closely with other related factors such as local language use and local terminology Wibowo et al. (2023). This survey brings out the need for situated studies of LLM-based applications in particular cultural contexts (e.g., restoring ancient texts from ancient cultures Assael et al. (2022); journalists in Africa Gondwe (2023)), which are conspicuously absent from the NLP literature. The combination of rigorous benchmarking and naturalistic studies will present a fuller picture of how culture plays out in LLMs.
@@
The survey is organized as follows. In Section 2, we describe our method for identifying the papers, categorizing them along various axes, and then deriving a taxonomy based on the proxies of cultures and probing methods used in the studies. These taxonomies are presented in Section 3 and Section 4 respectively, along with the papers that fall under various taxa. In Section 5, we discuss the gaps and recommendations, and finally, we conclude in Section 6.
2 Methodology
2.1 Identification and Categorization of Relevant Papers
@@
To compile a comprehensive list of papers that analyze the concept of “culture” within LLMs, our initial step involved an exhaustive search within the ACL Anthology [^4] https://aclanthology.org/ database with the keywords “culture” and “cultural”. We excluded research pertaining to speech recognition and multimodal systems, thereby narrowing down our survey to studies solely concerning text-based LLMs. We further searched on Google Scholar [^5] https://scholar.google.com/ for papers on “culture + LLM” and added the relevant papers to this list from conferences such as NeuRIPS [^6] https://neurips.cc and the Web Conference. [^7] https://www2024.thewebconf.org/ This initial search followed by a further manual filtering resulted in 39 papers published between 2020 and 2024.
@@
We read these papers and labeled them for (a) the definition of culture subscribed to in the paper, (b) the method used for probing the LLM for cultural awareness/bias, and (c) the languages and the cultures (thus defined) that were studied. It became apparent during the annotation process that none of the papers attempted to explicitly define “culture.” Therefore, instead of a definition, we decided to label the papers for (1) the proxy for culture used in a study and (2) the aspects of linguistic-culture Interaction that were studied. Using these labels, we then built taxonomies of the object of study and the method bottoms-up, which are described in this section, and also used to organize the survey findings in Sections 3 and 4.
@@ 2.2.1 Proxies of Culture
We identified 12 distinct labels for proxies of culture, which can further be categorized into two distinct groups:
1) Semantic Proxies: Often cultures are defined in terms of the emotions and values, food and drink (culinary cultures), kinship terms, social etiquette, etc. prevalent within a group of people. Thompson et al. (2020) groups these items under “semantic domains”, and they describe 21 semantic domains [^8] The complete list of semantic domains from Thompson et al. (2020) are: Quantity, time, kinship, function words, animals, sense perception, physical world, food and drink, cognition, possession, speech and language, spatial relations, the body, social and political relations, emotions and values, agriculture and vegetation, clothing and grooming, modern world, motion, basic actions and technology, the house. whose linguistic (and cognitive) usage is strongly influenced by culture. We use this framework to organize the semantic proxies of culture.
2) Demographic Proxies: As discussed earlier, culture is often described at the level of a community or group of people, who share certain common demographic attributes. These could be ethnicity (Masai culture), religion (Islamic culture), age (Gen Z culture), socio-economic class (middle class or urban), race, gender, language, region (Indonesian culture) and so on. Study of culture often describes communities at the intersection of these demographic proxies (e.g., Indian middle class).
Note that the semantic and demographic proxies are not orthogonal, for instance, one could define the moral values of a particular country as the cultural proxy to be studied.
2.2.2 Linguistic-Culture Interaction Axes
An alternative and useful approach to categorize the study of culture in language technology has been proposed by Hershcovich et al. (2022), where the interactions between language and culture is defined through three key dimensions:
Aboutness refers to the topics and issues that are prioritized or deemed relevant within different cultures. It notes how certain domains may be irrelevant or differently interpreted across cultures, affecting sentiment analysis and domain relevance in NLP. This aspect determines how cultural context influences what information is communicated. For instance, region-specific search logs on popular web search engines [^9] https://trends.google.com/trends/ show distinct patterns of interest which are partly influenced by local events and facts, but also partly due to differences in interests of the people, that is, the aboutness.
Common Ground is defined by the shared knowledge and assumptions among people within a culture. This aspect delves into conceptualization where languages and cultures shape the world into different semantic categories (e.g., kinship, spatial relations) and commonsense knowledge that represents a knowledge bank shared within a culture. Knowledge of local tradition, social norms and expected behaviors, dos and don’ts etc. are part of the cultural common ground.
Objectives and Values relate to the goals and ethical principles that guide behavior and communication. This aspect reflects the goals and ethical guidelines that inform the diversity in what different cultures value or prioritize. This dimension emphasizes the importance of aligning objectives with the ethical norms and values of various cultures, recognizing that these can significantly differ. Objectives and values discuss the ethical implications of NLP technologies, the need for fairness, and the challenges in balancing cultural preservation with minimizing biases.
@@ 2.3 Taxonomy: Probing Methods
There are two broad approaches to studying LLMs – the black-box approach which treats the LLM as a black-box and only relies on the observed responses to various inputs for analysis, and white-box approach where the internal states (such as the attention maps) of the models can be observed (e.g. Wichers et al. (2024)). Almost all studies we surveyed use the black-box approaches, where typically the input query is appended with a cultural conditioning or context and presented to the model. The responses of the model are compared under different cultural conditions as well as to baselines where no condition is present. These approaches can be further categorized as
@@
  • Generative Probing employs an open-ended fill-in-the-blank evaluation method for the LLMs and the text generated by the model under different cultural conditioning are compared.
@@ 3 Findings: Defining Culture
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Linguistic-Culture Interaction, xnode, l sep=2mm, [Objectives and Values, tnode, l sep = 2mm, [Kovač et al. (2023);Cao et al. (2023);Johnson et al. (2022);Dev et al. (2023);Gupta et al. (2024);Wan et al. (2023);Tanmay et al. (2023);An et al. (2023);Sorensen et al. (2023);Durmus et al. (2023);Cao et al. (2024a);, wnode]] [Common Ground, tnode, l sep = 2mm, [Koto et al. (2023);Wibowo et al. (2023);Wang et al. (2023);Johnson et al. (2022);Quan et al. (2020);Zhang et al. (2023);Bauer et al. (2023);Naous et al. (2023);Wu et al. (2023);Nguyen et al. (2023);Huang and Yang (2023a);Liu et al. (2024);,wnode] ] [Aboutness, tnode, l sep = 2mm, [Koto et al. (2024), wnode]]]
@@
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Demographic Proxies, xnode, l sep=2mm, [Region, tnode, l sep=2mm, [Hershcovich et al. (2022);Koto et al. (2023);Wibowo et al. (2023);Wang et al. (2023);Johnson et al. (2022);Wan et al. (2023);An et al. (2023);Zhang et al. (2023);Durmus et al. (2023);Jha et al. (2023);Ramezani and Xu (2023);Kabra et al. (2023);Zhou et al. (2023);Mukherjee et al. (2023);CH-Wang et al. (2023);Dev et al. (2023);Khanuja et al. (2023);Santy et al. (2023);Cao et al. (2023);Dwivedi et al. (2023);Koto et al. (2024);Cao et al. (2024a);Liu et al. (2024);,wnode]] [Language, tnode,l sep=2mm, [Hershcovich et al. (2022);Koto et al. (2023);Kovač et al. (2023);Wibowo et al. (2023);Cao et al. (2023);Johnson et al. (2022);Tanmay et al. (2023);Huang and Yang (2023a);Zhang et al. (2023);Kabra et al. (2023);Naous et al. (2023);Shaikh et al. (2023);Zhou et al. (2023);Mukherjee et al. (2023);CH-Wang et al. (2023);Dev et al. (2023);Khanuja et al. (2023);Santy et al. (2023);Das et al. (2023);Cao et al. (2024a);,wnode]] [Gender, tnode,l sep=2mm, [Cao et al. (2023);Johnson et al. (2022);Wan et al. (2023);An et al. (2023);Wu et al. (2023);,wnode]] [Race,tnode,l sep=2mm, [Johnson et al. (2022);Durmus et al. (2023);,wnode]] [Religion, tnode,l sep=2mm, [Koto et al. (2023);Durmus et al. (2023);Bauer et al. (2023);Das et al. (2023);,wnode]] [Education, tnode,l sep=2mm, [Koto et al. (2023);Quan et al. (2020);Bauer et al. (2023);Shaikh et al. (2023);Wu et al. (2023);Santy et al. (2023);,wnode]] [Ethnicity, tnode,l sep=2mm, [Koto et al. (2023);Johnson et al. (2022);Wan et al. (2023);Durmus et al. (2023);Santy et al. (2023);Cao et al. (2023);Das et al. (2023);Koto et al. (2024);Liu et al. (2024);,wnode] ]]
{forest}
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Semantic Proxies, xnode,l sep=2mm, [Names,tnode,l sep=2mm, [Wu et al. (2023);Quan et al. (2020);,wnode]] [Basic Actions and Technology,tnode,l sep=2mm, [Durmus et al. (2023);Bauer et al. (2023);,wnode]] [Social and Political Relations,tnode,l sep=2mm, [Wang et al. (2023);Johnson et al. (2022);Durmus et al. (2023);Shaikh et al. (2023);Bauer et al. (2023);Feng et al. (2023);Koto et al. (2024);,wnode]] [Food and Drink,tnode,l sep=2mm, [Palta and Rudinger (2023);Cao et al. (2024b);Koto et al. (2024);,wnode]] [Emotions and Values,tnode,l sep=2mm, [Hershcovich et al. (2022);Kovač et al. (2023);Koto et al. (2023);Wibowo et al. (2023);Cao et al. (2023);Johnson et al. (2022);Wan et al. (2023);Tanmay et al. (2023);Zhang et al. (2023);Shaikh et al. (2023);Jiang et al. (2022);Talat et al. (2021);Huang and Yang (2023b);Naous et al. (2023);Wu et al. (2023);Fung et al. (2023);Mukherjee et al. (2023);Santy et al. (2023);Cao et al. (2024b);Cao et al. (2024a);Liu et al. (2024);,wnode]] ]
@@
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Culture, root, l sep=4mm,s sep=10mm, [White-box Approach, fnode, l sep=3mm, [Mechanistic Interpretability, tnode, l sep = 3mm, [Wichers et al. (2024);, wnode] ] ] [Black-box Approach, fnode, l sep=3mm, [Discriminative Probing, tnode, l sep=3mm, [Cao et al. (2023);Tanmay et al. (2023);Rao et al. (2023);Kovač et al. (2023);,wnode]] [Generative Probing,tnode,l sep=3mm, [Nadeem et al. (2021);Nangia et al. (2020);Wan et al. (2023);Jha et al. (2023);,wnode]] ] ]

### v4

@@ Abstract
We present a survey of more than 90 recent papers that aim to study cultural representation and inclusion in large language models (LLMs). We observe that none of the studies explicitly define “culture”, which is a complex, multifaceted concept; instead, they probe the models on some specially designed datasets which represent certain aspects of “culture." We call these aspects the proxies of culture, and organize them across two dimensions of demographic and semantic proxies. We also categorize the probing methods employed. Our analysis indicates that only certain aspects of “culture,” such as values and objectives, have been studied, leaving several other interesting and important facets, especially the multitude of semantic domains Thompson et al. (2020) and aboutness Hershcovich et al. (2022), unexplored. Two other crucial gaps are the lack of robustness of probing techniques and situated studies on the impact of cultural mis- and under-representation in LLM-based applications.
@@
Muhammad Farid Adilazuarda1∗, Sagnik Mukherjee1thanks: Equal contribution, Pradhyumna Lavania2, Siddhant Singh2, Ashutosh Dwivedi2, Alham Fikri Aji1, Jacki O’Neill3, Ashutosh Modi2, Monojit Choudhury1 1MBZUAI  2Indian Institute of Technology Kanpur, India 3Microsoft Africa Research Institute, Kenya {farid.adilazuarda,sagnik.mukherjee}@mbzuai.ac.ae
@@ 1 Introduction
"Culture is the precipitate of cognition and communication in a human population." - Dan Sperber
Recently, there have been several studies on socio-cultural aspects of LLMs spanning from safety and value alignment Glaese et al. (2022); Bai et al. (2022b, a) to studying LLMs as personas belonging to certain cultures Gupta et al. (2024); Kovač et al. (2023) and their skills for resolving dilemmas in the context of value pluralism Sorensen et al. (2023); Tanmay et al. (2023).
@@
In order to make LLMs inclusive and deployable across regions and applications, it is indeed necessary for them to be able to function adequately under different “cultural” contexts. The growing body of work that broadly aims at evaluating LLMs for their multi-cultural awareness and biases underscore an important problem - that the existing models are strongly biased towards Western, Anglo-centric or American cultures Johnson et al. (2022); Cieciuch and Schwartz (2012); Dwivedi et al. (2023). Such biases are arguably detrimental to the performance of the models in non-Western contexts leading to disparate utility, potential for unfairness across regions. For instance, Haoyue and Cho (2024) and Chaves and Gerosa (2019) show that a conversational system that lacks cultural awareness alienate the users, leading to mistrust and lack of rapport, and eventual abandonment of the system by users from certain cultures. There are also concerns about the impact on global cultural diversity, since if biased models reinforce dominant cultures, whether implicitly or explicitly, they might lead to a cycle of cultural homogeneity Vaccino-Salvadore (2023); Schramowski et al. (2021). The recent generation of LLMs, with their impressive ability and widespread availability, only make this issue more pressing. It is therefore a timely moment to review the literature on LLMs and culture.
@@
In this work, we survey more than 90 NLP papers that study cultural representation, awareness or bias in LLMs either explicitly Huang and Yang (2023); Zhou et al. (2023b); Cao et al. (2024b) or implicitly Wan et al. (2023). It is quickly apparent that these papers either do not attempt to define culture or use very high-level definitions. For example, a common definition is “the way of life of a collective group of people, [that] distinguishes them from other groups with other cultures” (Mora, 2013; Shweder et al., 2007; Hershcovich et al., 2022). Not only do the papers typically use broad-brush definitions, most do not engage in a critical discussion on the topic. [^1] The situation is similar to that described in  Blodgett et al. (2020) in the context of research on “bias”. This is perhaps unsurprising as “culture” is a concept which evades simple definition.
1.1 Culture in the Social Sciences
Culture is multifaceted, meaning different things to different people at different times. For example, some of the many and often implicitly applied meanings of culture include: (a) “Cultural Heritage” such as art, music, and food habits [^2] https://uis.unesco.org/sites/default/files/documents/analysis_sdg_11.4.1_2022_final_alt_cover_0.pdf Blake (2000), (b) “Interpersonal Interactions” between people from different backgrounds (e.g., ways of speaking in a meeting, politeness norms) Monaghan et al. (2012), or (c) The “Ways of Life” of a collective group of people distinguishing them from other groups. There are a variety of sociological descriptions of culture, e.g., Parsons (1972) describes it as the the pattern of ideas and principles which abstractly specify how people should behave, but which do so in ways which prove practically effective relative to what people want to do (also see Münch et al. (1992)). However, these too are high-level and hard to concretise. Further complications arise because the instantiation of culture is necessarily situated. Every individual and group lies at the intersection of multiple cultures (defined by their political, professional, religious, regional, class-based and other affiliations) and these are invoked according to the situation, typically in contrast to another group(s).
In anthropology, a distinction has been made between thick and thin descriptions of culture Geertz (1973); Bourdieu (1972). Where culture as understood from the outsiders perspective, e.g. "people of type X believe in Y or behave in a particular manner" is a thin description of culture, as it does not consider the actor’s (of type X) personal perception of their context that resulted in that particular belief or the behavior. A thick description of culture, on the other hand, not only documents the observed behaviors but also the actors’ own explanations of the context and the behavior, and thus, can capture the insider-view of a culture as captured through people’s lived experiences.
1.2 Culture in NLP
How then is culture handled in NLP research? As we shall demonstrate, the datasets and studies are typically designed to tease out the differential performance of the models across some set of variables. Before we discuss these, we note that a couple of papers have begun to provide richer definitions of culture. Hershcovich et al. (2022) in their study calls out three axes of interaction between language and culture that NLP research and language technology needs to consider: common ground, aboutness and objectives and values. Aboutness refers to the topics and issues that are prioritized or deemed relevant within different cultures. Common Ground is defined by the shared knowledge and assumptions among people within a culture. Like the sociological and anthropological definitions of culture above, this provides a nice conceptualisation of culture, but practically it is hard to instantiate and measure in NLP studies. A recent survey paper Liu et al. (2024a) chooses a different definition of culture, based on White (1959) three dimensions of culture: 1) within human, 2) between humans, and 3) outside of human. Based on this, the paper creates a “taxonomy of culture" although the categorisation is a little complex.
In most of the NLP research seeking to examine culture, it is not defined at all beyond the high level. Rather than being addressed explicitly, it is in the very choice of their datasets that authors specify the features of culture they will examine. That is, the datasets themselves can be considered to be proxies for culture.
What do we mean by this? The authors of these papers investigating cultural representations in LLMs are seeking to understand how applicable LLMs are to different groups of people – and finding them apparently wanting in this count, they then seek to demonstrate and measure this concretely. Whilst they do not define culture beyond the high level (because, we would argue, a practical and actionable single definition of culture is hard to come by), the papers are still measuring some facet or other of cultural differences. The differences that they are measuring are instantiated in their datasets. For example, some papers examine food and drink, others differences in religious practices. These concrete, practical, measurable facets are in effect standing as proxies for culture. Since “cultures" are conceptual rather than concrete categories that are difficult to study directly through computational or quantitative methods, these proxies serve as easy to understand markers of culture that can be concretely captured through NLP datasets.
@@ 1.2 Culture in NLP
Given this wholly sensible strategy, it is useful to examine the different instantiations of culture found in this style of research. From food and drink, to norms and values, how have researchers represented culture in and through their datasets? In doing so we make explicit the various facets of culture which have been studied, and highlight gaps in the research. We call for a more explicit acknowledgment of the link between the datasets employed and the facets of culture studied, and hope that the schema described in this paper provides a useful mechanism for this.
@@
In addition, we highlight limitations in the robustness of the probing methods used in the studies, which raises doubts about the reliability and generalizability of the findings. Whilst benchmarking is important and necessary, it is not sufficient, as the choices made in creating rigorous benchmarking datasets are unlikely to reveal the full extent of either LLMs cultural limitations or their full cultural representation. Not only is culture multi-faceted, but cultural representation is tied in closely with other related factors such as local language use and local terminology Wibowo et al. (2023).
Our study also brings out the lack, and the urgent need thereof, for situated studies of LLM-based applications in particular cultural contexts (e.g., restoring ancient texts from ancient cultures Assael et al. (2022); journalists in Africa Gondwe (2023), and digital image making practices Mim et al. (2024)), which are conspicuously absent from the NLP literature. The combination of rigorous benchmarking and naturalistic studies will present a fuller picture of how culture plays out in LLMs.
@@
The survey is organized as follows. In Section 2, we describe our method for identifying the papers, categorizing them along various axes, and then deriving a taxonomy based on the proxies of cultures and probing methods used in the studies. These taxonomies are presented in Section 3 and Section 4 respectively. In Section 5, we discuss the gaps and recommendations. We conclude in Section 6.
2 Method
Scope of this survey is limited to the study of cultural representations within LLMs and LLM-based applications. Studies on culture in NLP that does not involve LLM have been excluded, and in order to keep this survey focused and manageable, we have also excluded studies on speech and multimodal models.
2.1 Searching Relevant Papers
@@
Our initial step is an exhaustive search within the ACL Anthology [^3] https://aclanthology.org/ database and a manual search on Google Scholar [^4] https://scholar.google.com/ for papers on culture and LLM, with the following keywords: “culture”, “cultural”,“culturally”, “norms”, “social”, “values”, “socio”, “moral”, “ethics”. We also searched for relevant papers from NeuRIPS [^5] https://neurips.cc and the Web Conference [^6] https://www2024.thewebconf.org/. This initial search followed by a manual filtering resulted in 90 papers published between 2020 and 2024.
@@
These papers were then manually labeled for (a) the definition of culture subscribed to in the paper, (b) the method used for probing the LLM for cultural awareness/bias, and (c) the languages and the cultures (thus defined) that were studied. It became apparent during the annotation process that none of the papers attempted to explicitly define “culture.” In the absense of definitions of culture, we labelled the papers according to (1) the types of data used to represent cultural differences which can be considered as a proxy for culture (as explained in Sec 1.2), and (2) the aspects of linguistic-culture interaction Hershcovich et al. (2022) that were studied. Using these labels, we then built taxonomies bottom-up for the object and the method of study.
@@ 2.2.1 Proxies of Culture
We identified 12 distinct labels into which the types of data or proxies of cultural difference can be categorized. These can be further classified into two overarching groups:
1) Demographic Proxies: Culture is, almost always, described at the level of a community or group of people, who share certain common demographic attributes. These could be ethnicity (Masai culture), religion (Islamic culture), age (Gen Z culture), socio-economic class (middle class or urban), race, gender, language, region (Indonesian culture) and so on, and their intersections (e.g., Indian middle class).
2) Semantic Proxies: Often cultures are defined in terms of the emotions and values, food and drink, kinship terms, social etiquette, etc. prevalent within a group of people. Thompson et al. (2020) groups these items under “semantic domains”, and they describe 21 semantic domains [^7] The complete list of semantic domains from Thompson et al. (2020) are: Quantity, time, kinship, function words, animals, sense perception, physical world, food and drink, cognition, possession, speech and language, spatial relations, the body, social and political relations, emotions and values, agriculture and vegetation, clothing and grooming, modern world, motion, basic actions and technology, the house. whose linguistic (and cognitive) usage is strongly influenced by culture. We use this framework to organize the semantic proxies of culture.
Note that the semantic and demographic proxies are orthogonal and simultaneously apply to any study. For instance one could choose to study the festivals (a semantic proxy) celebrated in a particular country (a demographic proxy).
@@ 2.3 Taxonomy: Probing Methods
There are two broad approaches to studying LLMs – the black-box approach which treats the LLM as a black-box and only relies on the observed responses to various inputs for analysis, and white-box approach where the internal states (such as the attention maps) of the models can be observed e.g. Wichers et al. (2024). Almost all studies we surveyed use the black-box approaches, where typically the input query is appended with a cultural context and presented to the model. The responses of the model are compared under different cultural conditions as well as to baselines where no condition is present. These approaches can be further categorized as
@@
  • Generative Probing uses an open-ended fill-in-the-blank evaluation method for the LLMs and the text generated by the model under different cultural conditioning are compared.
@@ 3 Findings: Defining Culture
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Demographic Proxies, xnode, l sep=2mm, [Region, tnode, l sep=2mm, [Koto et al. (2023);Wibowo et al. (2023);Wang et al. (2023);Johnson et al. (2022);Wan et al. (2023);An et al. (2023);Zhang et al. (2023);Durmus et al. (2023);Jha et al. (2023);Ramezani and Xu (2023);Zhou et al. (2023b);Mukherjee et al. (2023);CH-Wang et al. (2023);Dev et al. (2023);Khanuja et al. (2023);Santy et al. (2023);Cao et al. (2023);Dwivedi et al. (2023);Koto et al. (2024);Cao et al. (2024a);Liu et al. (2024b);Masoud et al. (2024);Nguyen et al. (2024);Lee et al. (2023);Zhou et al. (2023a);Chiu et al. (2024);Atari et al. (Working Paper),wnode]] [Language, tnode,l sep=2mm, [Koto et al. (2023);Kovač et al. (2023);Cao et al. (2023);Cao et al. (2023);Johnson et al. (2022);Huang and Yang (2023);Zhang et al. (2023);Kabra et al. (2023);Naous et al. (2023);Shaikh et al. (2023);Zhou et al. (2023b);Mukherjee et al. (2023);CH-Wang et al. (2023);Dev et al. (2023);Khanuja et al. (2023);Santy et al. (2023);Das et al. (2023);Cao et al. (2024a);Havaldar et al. (2023);Mohamed et al. (2022);akinade-etal-202-varepsilon; Ventura et al. (2023); Buttrick (2024);Luo et al. (2024); Choenni et al. (2024); Keleg and Magdy (2023),wnode]] [Gender, tnode,l sep=2mm, [Johnson et al. (2022);Wan et al. (2023);Wu et al. (2023);Frenda et al. (2023);Caliskan et al. (2017),wnode]] [Race,tnode,l sep=2mm, [Johnson et al. (2022);Durmus et al. (2023);Hwang et al. (2023);Pei and Jurgens (2023);Durmus et al. (2024);Cooper et al. (2024),wnode]] [Religion, tnode,l sep=2mm, [Koto et al. (2023);Durmus et al. (2023);Bauer et al. (2023);Das et al. (2023);Nguyen et al. (2023);Li et al. (2024b);Durmus et al. (2024);Keleg and Magdy (2023),wnode]] [Education, tnode,l sep=2mm, [Koto et al. (2023);Quan et al. (2020);Bauer et al. (2023);Wu et al. (2023);Santy et al. (2023);Zhao et al. (2024);AlKhamissi et al. (2024);Hwang et al. (2023);Beck et al. (2024a);Li et al. (2024b);Son et al. (2024);Kirk et al. (2024);Kim et al. (2024); Chiu et al. (2024),wnode]] [Ethnicity, tnode,l sep=2mm, [Koto et al. (2023);Johnson et al. (2022);Wan et al. (2023);Durmus et al. (2023);Santy et al. (2023);Koto et al. (2024);Sap et al. (2022);Shi et al. (2024);Durmus et al. (2024);Cooper et al. (2024);Kirk et al. (2024); Chiu et al. (2024),wnode] ]]
@@
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Semantic Proxies, xnode,l sep=2mm, [Names,tnode,l sep=2mm, [Aher et al. (2023);Rai et al. (2024);Sandoval et al. (2023),wnode]] [Basic Actions and Technology,tnode,l sep=2mm, [Durmus et al. (2023);Zhao et al. (2024);Zhan et al. (2023);Zhan et al. (2024);Bhatia and Shwartz (2023);Ringel et al. (2019);Choenni et al. (2024);Ziems et al. (2023),wnode]] [Social and Political Relations,tnode,l sep=2mm, [Johnson et al. (2022);Durmus et al. (2023);Shaikh et al. (2023);Feng et al. (2023);Koto et al. (2024);Forbes et al. (2020);Masoud et al. (2024);Beck et al. (2024a);Li et al. (2024b);Santurkar et al. (2023)Li et al. (2024a);Lee et al. (2023);Cooper et al. (2024);Ziems et al. (2023);Jin et al. (2024);Kim et al. (2024),wnode]] [Food and Drink,tnode,l sep=2mm, [Palta and Rudinger (2023);Cao et al. (2024b);Koto et al. (2024);Fung et al. (2024);Nguyen et al. (2023);Yao et al. (2024);Putri et al. (2024);Li et al. (2024b);Zhou et al. (2024);Kirk et al. (2024),wnode]] [Emotions and Values,tnode,l sep=2mm, [Hershcovich et al. (2022);Kovač et al. (2023);Koto et al. (2023);Wibowo et al. (2023);Cao et al. (2023);Johnson et al. (2022);Wan et al. (2023);Tanmay et al. (2023);Zhang et al. (2023);Shaikh et al. (2023);Jiang et al. (2022);Talat et al. (2021);Huang and Yang (2023);Naous et al. (2023);Wu et al. (2023);Fung et al. (2023);Mukherjee et al. (2023);Santy et al. (2023);Cao et al. (2024b);Cao et al. (2024a);Liu et al. (2024b);Friedrich et al. (2023);Havaldar et al. (2023);Moghimifar et al. (2023);Rao et al. (2023b),wnode]] ]
@@
for tree= grow=east, growth parent anchor=west, parent anchor=east, child anchor=west, calign = edge midpoint, [Probing methods, root, l sep=4mm,s sep=10mm, [White-box Approach, fnode, l sep=3mm, [Mechanistic Interpretability, tnode, l sep = 3mm, [Wichers et al. (2024);, wnode] ] ] [Black-box Approach, fnode, l sep=3mm, [Discriminative Probing, tnode, l sep=3mm, [Cao et al. (2023);Tanmay et al. (2023);Rao et al. (2023a);Kovač et al. (2023);,wnode]] [Generative Probing,tnode,l sep=3mm, [Nadeem et al. (2021);Nangia et al. (2020);Wan et al. (2023);Jha et al. (2023);Li et al. (2024c);,wnode]] ] ]